How many different four letter english words(with or without meaning) can be written using distinct letters from alphabet?


1
Expert's answer
2021-03-31T01:58:01-0400

The alphabet has 26 letters. We can write "26\cdot 25\cdot 24\cdot 23=3588000" different words. We used multiplication principle from combinatorics.